# Missing Pieces: How Framing Uncertainty Impacts Longitudinal Trust in AI Decision Aids — A Gig Driver Case Study

### v1

@@ 8. Conclusion
In this work, we investigated how different framings of uncertain estimates impacted the extent to which users were willing to trust in and rely on an AI decision aid within a context characterised by uncertainty and repeated interactions. As a case study, we deployed these framings as part of a longitudinal trial of a schedule recommendation tool to 51 gig drivers. Our quantitative results affirm the influence of users’ perceptions of the AI’s accuracy in estimating outcomes on trust and reliance, and demonstrate that framings of uncertainty have varying impacts on trust and reliance. While increased granularity and hedging language did not appear to improve trust like ranged estimates did, they did lead to participants expressing higher reliance and exhibiting lower attrition.
The most prominent conclusion we drew from the interviews we conducted, however, is that there is no one-size-fits-all approach to building trust. Even within the same condition, participants exhibited great variability in how they incorporated the AI system’s recommendations into their planning, and in how they reacted to the outcomes of their reliance. We thus echo calls to adopt user-centred design approaches for deployments of AI systems (Rantavuo, 2019; Wang et al., 2019). In particular, it is not enough to consider users as a homogeneous group. Designing trustworthy AI decision aids requires recognising the diverse needs and perceptions of different groups of users, and aligning explanations to the expectations of each group. Ideally, future empirical work should control more precisely for different user factors to support the identification of these groups. However, we also recognise that this is likely to be difficult: longitudinal effects such as those we uncovered imply that users’ latent attitudes cannot be easily assessed in a prospective manner, and users’ stated and revealed attitudes may also differ (Fietta et al., 2021).
Among the characteristics of users that are likely to shape their trust in a specific AI decision aid is their innate propensity to trust AI systems in general (Solberg et al., 2022). Various empirical studies have found users to exhibit either propensities towards algorithmic aversion (Mahmud et al., 2022) — a reluctance to rely on algorithms — or its polar opposite, algorithmic appreciation (Logg et al., 2019). These results are less likely to be contradictory than reflective of genuine attitudes that exist in different segments of the population; (Stein et al., 2024), for instance, found that algorithmic appreciation correlated with personality traits and age. Our longitudinal results support findings in single-shot settings that the context-specific expression of these general attitudes is influenced by framing (Hou and Jung, 2021) and outcomes (Dietvorst et al., 2015). But to what extent do specific interactions with AI systems conversely influence users’ general attitudes? We believe that this question merits further exploration.
More fundamentally, what should an AI decision aid’s goal be in the presence of uncertainty? Should it be to ensure the user’s compliance or simply their reliance? We demonstrate that reliance and compliance are not always aligned, and therefore that a trade-off must be made in certain contexts. However, we concur with (Schemmer et al., 2023) in that, if the nominal purpose of an AI decision aid is to support the user’s own goals, then a notion of appropriate reliance should take precedence over strict compliance due to the potential for misalignment in objectives. Since both the behaviour (Carroll et al., 2022) and presentation (Vasconcelos et al., 2023) of AI systems can both impact the reliance-compliance trade-off in users, we believe that careful requirements engineering is necessary to design AI decision aids with practical utility.
We viewed gig driving as a representative context that could provide rich data for our study. However, the challenges of uncertainty and repeated interactions are also prevalent in other deployment contexts of AI decision aids, such as healthcare, education, and criminal justice. We believe that our findings are sufficiently general that they could readily apply to other contexts; yet, the idiosyncrasies of different contexts may also influence the exact mechanisms by which framings of uncertainty impact trust and reliance. For instance, unlike how gig drivers receive near-instant feedback through earnings in our context, medical practitioners may not observe the impacts of their decisions until long afterwards. Our study is far from being the final word on how the framing of uncertainty impacts trust, but it suggests a fruitful line of future work tailored towards different deployment contexts of AI systems.
@@ Acknowledgements.
We thank Joshua Mateer, Jenny T. Liang, Jose M. del Alamo, Ian Yang, Chenyang Yang, Stephanie Milani, and Niklas Kühl for their invaluable feedback and discussion on the study’s instruments, protocol, analysis, and manuscript. Also, we thank Linda Moreci for assistance with participant compensation, and Akshara Khare and Anagha Ravi Shankara for their contributions to the recommendation tool’s implementation. This work was conducted in collaboration with Gridwise; we thank Kunal Prajapati, Karan Moudgil, Ryan Green, and Brian Finnamore for their assistance with recruitment and data procurement. We were also supported by a research grant from Mobility21 (a US DOT National University Transportation Center) and the Tang Family Endowed Innovation Fund.References
@@
- Aoki (2021) Naomi Aoki. 2021. The importance of the assurance that “humans are still in the decision loop” for public trust in artificial intelligence: Evidence from an online experiment. Computers in Human Behavior 114 (2021), 106572.
- Bansal et al. (2019) Gagan Bansal, Besmira Nushi, Ece Kamar, Walter S. Lasecki, Daniel S. Weld, and Eric Horvitz. 2019. Beyond Accuracy: The Role of Mental Models in Human-AI Team Performance. In Proceedings of the 7th AAAI Conference on Human Computation and Crowdsourcing (HCOMP ’19). AAAI, Stevenson, USA, 2–11.
@@
- Cai et al. (2019a) Carrie J. Cai, Samantha Winter, David Steiner, Lauren Wilcox, and Michael Terry. 2019a. “Hello AI”: Uncovering the Onboarding Needs of Medical Practitioners for Human–AI Collaborative Decision-Making. Proceedings of the ACM on Human-Computer Interaction 3 (2019), 1–24.
- Cai et al. (2019b) Carrie J. Cai, Samantha Winter, David Steiner, Lauren Wilcox, and Michael Terry. 2019b. Reliability and Inter-rater Reliability in Qualitative Research: Norms and Guidelines for CSCW and HCI Practice. Proceedings of the ACM on Human-Computer Interaction 3 (2019), 1–23.
- Cameron et al. (2022) Chris Cameron, Jason Hartford, Taylor Lundy, and Kevin Leyton-Brown. 2022. The Perils of Learning Before Optimizing. In Proceedings of the 36th AAAI Conference on Artificial Intelligence (AAAI ’21). AAAI, Virtual, 3708–3715.
- Carroll et al. (2022) Micah D Carroll, Anca Dragan, Stuart Russell, and Dylan Hadfield-Menell. 2022. Estimating and Penalizing Induced Preference Shifts in Recommender Systems. In Proceedings of the 39th International Conference on Machine Learning (ICML ’22). PMLR, Hawaii, 2686–2709.
@@
- Cavanaugh and Neath (2019) Joseph E. Cavanaugh and Andrew A. Neath. 2019. The Akaike information criterion: Background, derivation, properties, application, interpretation, and refinements. Wiley Interdisciplinary Reviews: Computational Statistics 11, 3 (2019), e1460.
- Chen et al. (2023) Yiqun T. Chen, Katharina Reinecke, Angela D. R. Smith, and Alexandra To. 2023. Why, when, and from whom: considerations for collecting and reporting race and ethnicity data in HCI. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Hamburg, Germany, 1–15.
@@
- Dietvorst et al. (2015) Berkeley J. Dietvorst, Joseph P. Simmons, and Cade Massey. 2015. Algorithm aversion: People erroneously avoid algorithms after seeing them err. Journal of Experimental Psychology: General 144, 1 (2015), 114–126.
@@
- Ferraro et al. (2020) Andres Ferraro, Dietmar Jannach, and Xavier Serra. 2020. Exploring Longitudinal Effects of Session-based Recommendations. In Proceedings of the 14th ACM Conference on Recommender Systems (RecSys ’20). ACM, Virtual, 474–479.
- Fietta et al. (2021) Valentina Fietta, Francesca Zecchinato, Brigida Di Stasi, Mirko Polato, and Merylin Monaro. 2021. Dissociation Between Users’ Explicit and Implicit Attitudes Toward Artificial Intelligence: An Experimental Study. IEEE Transactions on Human-Machine Systems 52, 3 (2021), 481–489.
@@
- Gillath et al. (2021) Omri Gillath, Ting Ai, Michael S. Branicky, Shawn Keshmiri, Robert B. Davison, and Ryan Spaulding. 2021. Attachment and trust in artificial intelligence. Computers in Human Behavior 115 (2021), 106607.
- Gretzel and Fesenmaier (2006) Ulrike Gretzel and Daniel R. Fesenmaier. 2006. Persuasion in Recommender Systems. International Journal of Electronic Commerce 11, 2 (2006), 81–100.
@@
- Ho et al. (2013) Geoffrey Ho, Nada Pavlovic, Vincent Myers, and Robert Arrabito. 2013. Reducing False Alarms in Automated Target Recognition by Lowering the Level of Automation. In Proceedings of the 2013 Human Factors and Ergonomics Society Annual Meeting (HFES ’13). SAGE, San Diego, USA, 374–378.
- Holmström and Hällgren (2022) Jonny Holmström and Markus Hällgren. 2022. AI management beyond the hype: exploring the co-constitution of AI and organizational context. AI & Society 37 (2022), 1575–1585.
- Hou and Jung (2021) Yoyo Tsung-Yu Hou and Malte F. Jung. 2021. Who is the Expert? Reconciling Algorithm Aversion and Algorithm Appreciation in AI-Supported Decision Making. Proceedings of the ACM on Human-Computer Interaction 5 (2021), 1–25.
@@
- Jiang et al. (2004) Xiaochun Jiang, Mohammad T. Khasawneh, Reena Master, Shannon R. Bowling, Anand K. Gramopadhye, Brian J. Melloy, and Larry Grimes. 2004. Measurement of human trust in a hybrid inspection system based on signal detection theory measures. International Journal of Industrial Ergonomics 34, 5 (2004), 407–419.
- Kaplan et al. (2023) Alexandra D. Kaplan, Theresa T. Kessler, J. Christopher Brill, and P. A. Hancock. 2023. Trust in Artificial Intelligence: Meta-Analytic Findings. Human Factors 65, 2 (2023), 337–359.
- Kaufman and Kirsh (2022) Robert A. Kaufman and David J. Kirsh. 2022. Cognitive Differences in Human and AI Explanation. In Proceedings of the 44th Annual Conference of the Cognitive Science Society (CogSci ’22). eScholarship, Toronto, Canada, 2694–2700.
@@
- Komiak and Benbasat (2006) Sherrie Y. X. Komiak and Izak Benbasat. 2006. The Effects of Personalization and Familiarity on Trust and Adoption of Recommendation Agents. MIS Quarterly 30, 4 (2006), 941–960.
@@
- Lee and Moray (1992) John Lee and Neville Moray. 1992. Trust, control strategies and allocation of function in human-machine systems. Ergonomics 10 (1992), 1243–1270.
- Lee and See (2004) John D. Lee and Katrina A. See. 2004. Trust in Automation: Designing for Appropriate Reliance. Human Factors 46, 1 (2004), 50–80.
@@
- Liang and Willemsen (2022) Yu Liang and Martijn C. Willemsen. 2022. Exploring the longitudinal effects of nudging on users’ music genre exploration behavior and listening preferences. In Proceedings of the 16th ACM Conference on Recommender Systems (RecSys ’22). ACM, Seattle, USA, 3–13.
- Liao and Sundar (2022a) Mengqi Liao and S. Shyam Sundar. 2022a. When E-Commerce Personalization Systems Show and Tell: Investigating the Relative Persuasive Appeal of Content-Based versus Collaborative Filtering. Journal of Advertising 51 (2022), 256–267.
- Liao and Sundar (2022b) Q. Vera Liao and S. Shyam Sundar. 2022b. Designing for Responsible Trust in AI Systems: A Communication Perspective. In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’22). ACM, Seoul, Korea, 1257–1268.
- Lockey et al. (2021) Steven Lockey, Nicole Gillespie, Daniel Holm, and Ida Asadi Someh. 2021. A Review of Trust in Artificial Intelligence: Challenges, Vulnerabilities and Future Directions. In Proceedings of the 54th Hawaii International Conference on System Sciences (HICSS ’21). AIS, Honolulu, USA, 5463–5472.
- Logg et al. (2019) Jennifer M. Logg, Julia A. Minson, and Don A. Moore. 2019. Algorithm appreciation: People prefer algorithmic to human judgment. Organizational Behavior and Human Decision Processes 151 (2019), 90–103.
@@
- Mahmud et al. (2022) Hasan Mahmud, A.K.M. Najmul Islam, Syed Ishtiaque Ahmed, and Kari Smolander. 2022. What influences algorithmic decision-making? A systematic literature review on algorithm aversion. Technological Forecasting and Social Change 175 (2022), 121390.
@@
- McGregor (2021) Sean McGregor. 2021. Preventing Repeated Real World AI Failures by Cataloging Incidents: The AI Incident Database. In Proceedings of the 35th AAAI Conference on Artificial Intelligence (AAAI ’21). AAAI, Virtual, 15458–15463.
- Mcknight et al. (2011) D. Harrison Mcknight, Michelle Carter, Jason Bennett Thatcher, and Paul F. Clay. 2011. Trust in a specific technology: An investigation of its components and measures. ACM Transactions on Management Information Systems 2, 2 (2011), 1–25.
- Meske and Bunde (2020) Christian Meske and Enrico Bunde. 2020. Transparency and Trust in Human-AI-Interaction: The Role of Model-Agnostic Explanations in Computer Vision-Based Decision Support. In Proceedings of the 2020 International Conference on Human-Computer Interaction (HCII ’20). Springer, Copenhagen, Denmark, 54–69.
@@
- Naiseh et al. (2021) Mohammad Naiseh, Reem S. Al-Mansoori, Dena Al-Thani, Nan Jiang, and Raian Ali. 2021. Nudging through Friction: An Approach for Calibrating Trust in Explainable AI. In Proceedings of the 8th International Conference on Behavioral and Social Computing (BESC ’21). IEEE, Virtual, 1–5.
@@
- Nourani et al. (2021) Mahsan Nourani, Chiradeep Roy, Jeremy E. Block, Donald R. Honeycutt, Tahrima Rahman, Eric D. Ragan, and Vibhav Gogate. 2021. Anchoring Bias Affects Mental Model Formation and User Reliance in Explainable AI Systems. In Proceedings of the 26th International Conference on Intelligent User Interfaces (IUI ’21). ACM, College Station, USA, 340–350.
@@
- Rader et al. (2018) Emilee Rader, Kelley Cotter, and Janghee Cho. 2018. Explanations as Mechanisms for Supporting Algorithmic Transparency. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (CHI ’18). ACM, Montréal, Canada, 1–14.
- Rantavuo (2019) Heli Rantavuo. 2019. Designing for intelligence: User-centred design in the age of algorithms. In Proceedings of the 5th International ACM In-Cooperation HCI and UX Conference (CHIuXiD ’23). ACM, Bali, Indonesia, 182–187.
@@
- Saßmannshausen et al. (2021) Till Saßmannshausen, Peter Burggräf, Johannes Wagner, Marc Hassenzahl, Thomas Heupel, and Fabian Steinberg. 2021. Trust in artificial intelligence within production management — an exploration of antecedents. Ergonomics 64, 10 (2021), 1333–1350.
@@
- Shin (2021) Donghee Shin. 2021. The effects of explainability and causability on perception, trust, and acceptance: Implications for explainable AI. International Journal of Human-Computer Studies 146 (2021), 102551.
@@
- Stein et al. (2023) Jake M L Stein, Vidminas Vizgirda, Max Van Kleek, Reuben Binns, Jun Zhao, Rui Zhao, Naman Goel, George Chalhoub, Wael S Albayaydh, and Nigel Shadbolt. 2023. ‘You are you and the app. There’s nobody else.’: Building Worker-Designed Data Institutions within Platform Hegemony. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Hamburg, Germany, 1–26.
- Stein et al. (2024) Jan-Philipp Stein, Tanja Messingschlager, Timo Gnambs, Fabian Hutmacher, and Markus Appel. 2024. Attitudes towards AI: measurement and associations with personality. Scientific Reports 14 (2024), 2909.
@@
- Sun et al. (2023) Yuan Sun, Magdalayna Drivas, Mengqi Liao, and S. Shyam Sundar. 2023. When Recommender Systems Snoop into Social Media, Users Trust them Less for Health Advice. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Hamburg, Germany, 1–14.
- Sundar et al. (2008) S. Shyam Sundar, Anne Oeldorf-Hirsch, and Qian Xu. 2008. The Bandwagon Effect of Collaborative Filtering Technology. In Proceedings of the 2008 CHI Conference on Human Factors in Computing Systems (CHI ’08). ACM, Florence, Italy, 3453–3458.
- Suwanaposee et al. (2023) Pang Suwanaposee, Carl Gutwin, Zhe Chen, and Andy Cockburn. 2023. ‘Specially For You’ — Examining the Barnum Effect’s Influence on the Perceived Quality of System Recommendations. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). ACM, Hamburg, Germany, 1–11.
@@
- Textor et al. (2022) Claire Textor, Rui Zhang, Jeremy Lopez, Beau G. Schelble, Nathan J. McNeese, Guo Freeman, Richard Pak, Chad Tossell, and Ewart J. de Visser. 2022. Exploring the Relationship Between Ethics and Trust in Human–Artificial Intelligence Teaming: A Mixed Methods Approach. Journal of Cognitive Engineering and Decision Making 16, 4 (2022), 252–281.
@@
- Vasconcelos et al. (2023) Helena Vasconcelos, Matthew Jörke, Madeleine Grunde-McLaughlin, Tobias Gerstenberg, Michael S. Bernstein, and Ranjay Krishna. 2023. Explanations Can Reduce Overreliance on AI Systems During Decision-Making. Proceedings of the ACM on Human-Computer Interaction 7 (2023), 1–38.
@@
- Wang et al. (2019) Danding Wang, Qian Yang, Ashraf Abdul, and Brian Y. Lim. 2019. Designing Theory-Driven User-Centric Explainable AI. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (CHI ’19). ACM, Glasgow, UK, 1–15.
@@
- Wintersberger et al. (2021) Philipp Wintersberger, Frederica Janotta, Jakob Peintner, Andreas Löcken, and Andreas Riener. 2021. Evaluating feedback requirements for trust calibration in automated vehicles. it — Information Technology 63, 2 (2021), 1–12.
@@
- Wu et al. (2019) Qingjun Wu, Hao Zhang, Zhen Li, and Kai Liu. 2019. Labor control in the gig economy: Evidence from Uber in China. Journal of Industrial Relations 61, 4 (2019), 574–596.
@@
- Zhang et al. (2022) Angie Zhang, Alexander Boltz, Chun-Wei Wang, and Min Kyung Lee. 2022. Algorithmic Management Reimagined For Workers and By Workers: Centering Worker Well-Being in Gig Work. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). ACM, New Orleans, USA, 1–20.
@@
- Zhou et al. (2021) Meizi Zhou, Jingjng Zhang, and Gediminas Adomavicius. 2021. Longitudinal Impact of Preference Biases on Recommender Systems’ Performance. Technical Report 2021-10. Kelley School of Business Research Paper.

### v2

@@ 8. Conclusion
In this paper, we assessed how users’ trust and reliance on AI decision aids is influenced by designs that expose uncertainty. Unlike the laboratory experiments used by previous work, we did so within the real-world context of gig driving. Specifically, we conducted a longitudinal, in situ user study of an AI-based schedule recommendation tool with gig drivers. These drivers’ interactions with our tool impacted their actual earnings. Our findings demonstrate that trust can be built by (1) maintaining perceptions of accuracy over repeated interactions and (2) displaying uncertainty in a task-aligned fashion, the latter of which points to a need for more context-specific evaluations of AI decision aids.
@@ Acknowledgements.
We thank Jenny T. Liang, Joshua Mateer, Jose M. del Alamo, Ian Yang, Chenyang Yang, Kaia Newman, Stephanie Milani, Niklas Kühl, and Nadia Nahar for their feedback and discussion on the study’s instruments, protocol, analysis, and manuscript. Also, we thank Linda Moreci for assistance with participant compensation, and Akshara Khare and Anagha Ravi Shankara for contributions to the recommendation tool’s implementation. This work was conducted in collaboration with Gridwise; we thank Kunal Prajapati, Karan Moudgil, Ryan Green, and Brian Finnamore for help with recruitment and data procurement. We were supported by a research grant from Mobility21 and the Tang Family Endowed Innovation Fund.References
@@
- Bates et al. (2015) Douglas Bates, Martin Mächler, Ben Bolker, and Steve Walker. 2015. Fitting Linear Mixed-Effects Models Using lme4. Journal of Statistical Software 67, 1 (2015), 1–48.
@@
- Cai et al. (2019) Carrie J. Cai, Samantha Winter, David Steiner, Lauren Wilcox, and Michael Terry. 2019. “Hello AI”: Uncovering the Onboarding Needs of Medical Practitioners for Human–AI Collaborative Decision-Making. Proceedings of the ACM on Human-Computer Interaction 3 (2019), 1–24.
- Cai et al. (2021) Carrie J. Cai, Samantha Winter, David Steiner, Lauren Wilcox, and Michael Terry. 2021. Onboarding Materials as Cross-functional Boundary Objects for Developing AI Assistants. In Extended Abstracts of the 2021 CHI Conference on Human Factors in Computing Systems (CHI EA ’21). ACM, Yokohama, Japan, 1–7.
@@
- Jacovi et al. (2021) Alon Jacovi, Ana Marasović, Tim Miller, and Yoav Goldberg. 2021. Formalizing Trust in Artificial Intelligence: Prerequisites, Causes and Goals of Human Trust in AI. In Proceedings of the 2021 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’21). ACM, Chicago, USA, 624–635.
@@
- Kim et al. (2024) Sunnie S. Y. Kim, Q. Vera Liao, Mihaela Vorvoreanu, Stephanie Ballard, and Jennifer Wortman Vaughan. 2024. “I’m Not Sure, But…”: Examining the Impact of Large Language Models’ Uncertainty Expression on User Reliance and Trust. In Proceedings of the 2024 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’24). ACM, Rio de Janeiro, Brazil, 822–835.
- Kim et al. (2023) Sunnie S. Y. Kim, Elizabeth Anne Watkins, Olga Russakovsky, Ruth Fong, and Andrés Monroy-Hernández. 2023. Humans, AI, and Context: Understanding End-Users’ Trust in a Real-World Computer Vision Application. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’23). ACM, Chicago, USA, 77–88.
- Knowles and Richards (2021) Bran Knowles and John T. Richards. 2021. The Sanction of Authority: Promoting Public Trust in AI. In Proceedings of the 2021 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’21). ACM, Chicago, USA, 262–271.
@@
- Kuznetsova et al. (2017) Alexandra Kuznetsova, Per B. Brockhoff, and Rune H. B. Christensen. 2017. lmerTest Package: Tests in Linear Mixed Effects Models. Journal of Statistical Software 82, 13 (2017), 1–26.
@@
- Liao and Sundar (2022) Q. Vera Liao and S. Shyam Sundar. 2022. Designing for Responsible Trust in AI Systems: A Communication Perspective. In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’22). ACM, Seoul, Korea, 1257–1268.
@@
- Macrae (2021) Carl Macrae. 2021. Learning from the Failure of Autonomous and Intelligent Systems: Accidents, Safety, and Sociotechnical Sources of Risk. Risk Analysis 42, 9 (2021), 1999–2025.
@@
- Pareek et al. (2024) Saumya Pareek, Eduardo Velloso, and Jorge Goncalves. 2024. Trust Development and Repair in AI-Assisted Decision-Making during Complementary Expertise. In Proceedings of the 2024 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’24). ACM, Rio de Janeiro, Brazil, 546–561.
@@
- Stanley and Dorton (2023) Jeff C. Stanley and Stephen L. Dorton. 2023. Exploring Trust With the AI Incident Database. In Proceedings of the 2023 Human Factors and Ergonomics Society Annual Meeting (HFES ’23). ACM, Washington DC, USA, 489–494.
@@
- Widder et al. (2021) David Gray Widder, Laura Dabbish, James D. Herbsleb, Alexandra Holloway, and Scott Davidoff. 2021. Trust in Collaborative Automation in High Stakes Software Engineering Work: A Case Study at NASA. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). ACM, Yokohama, Japan, 1–13.
@@
- Wysocki et al. (2023) Oskar Wysocki, Jessica Katharine Davies, Markel Vigo, Anne Caroline Armstrong, Dónal Landers, Rebecca Lee, and André Freitas. 2023. Assessing the communication gap between AI models and healthcare professionals: Explainability, utility and trust in AI-driven clinical decision-making. Artificial Intelligence 316 (2023), 103839.
- Yang et al. (2024) Fumeng Yang, Chloe Rose Mortenson, Erik Nisbet, Nicholas Diakopoulos, and Matthew Kay. 2024. In Dice We Trust: Uncertainty Displays for Maintaining Trust in Election Forecasts Over Time. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (CHI ’24). ACM, Honolulu, USA, 1–24.
@@
- Yu et al. (2019) Kun Yu, Shlomo Berkovsky, Ronnie Taib, Jianlong Zhou, and Fang Chen. 2019. Do I trust my machine teammate? An investigation from perception to decision. In Proceedings of the 24th International Conference on Intelligent User Interfaces (IUI ’19). ACM, Los Angeles, USA, 460–468.
@@
- Zhang et al. (2022a) Angie Zhang, Alexander Boltz, Chun-Wei Wang, and Min Kyung Lee. 2022a. Algorithmic Management Reimagined For Workers and By Workers: Centering Worker Well-Being in Gig Work. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). ACM, New Orleans, USA, 1–20.
- Zhang et al. (2024) Dongping Zhang, Jason Hartline, and Jessica Hullman. 2024. Designing Shared Information Displays for Agents of Varying Strategic Sophistication. In Proceedings of the 27st ACM Conference on Computer-Supported Cooperative Work and Social Computing (CSCW ’24). ACM, San José, Costa Rica, 1–34.
- Zhang et al. (2022b) Qiaoning Zhang, Matthew L. Lee, and Scott Carter. 2022b. You Complete Me: Human-AI Teams and Complementary Expertise. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). ACM, New Orleans, USA, 1–28.